# Pyrometheus: Symbolic abstractions for XPU and automatically differentiated computation of combustion kinetics and thermodynamics

### v1

@@ 1.1 Computational Challenges of Combustion Thermochemistry
The evaluation of the thermochemistry often adds significantly to the cost of simulations of reacting flows. This cost primarily follows three sources. The first is an increase in the number of transported variables with respect to simulations of inert flow; the mixture composition can be described by up to hundreds of reacting scalars. The second source is the evaluation of up to thousands of exponentials that appear in expressions representing reaction rates. Last is a time-step restriction that increases the time of solution and stems from the imposed chemical time scale, over 1000 times smaller than that of the flow time scale [3]. These effect compound and reacting flow simulations can be prohibitively expensive. Exascale computers can mitigate this cost by using compute accelerators (XPUs), such as graphic processing units (GPUs) and accelerated processing units (APUs, or superchips).
Our approach complements established thermochemistry software, such as dedicated libraries like Cantera [4], Chemkin [5], and TChem [6], which provide broad modeling capabilities but are limited to CPU hardware. While serving the community well, these are not easily optimized for high-performance simulation, especially on accelerator hardware. This is because the compiled libraries are standalone and live in a different executable than the flow solver. Because of their breadth and structure, adapting such libraries to GPUs is an extensive task that risks conflicting with the goals and success of these tools. Writing new thermochemistry software exclusively for large-scale flow simulations on accelerators risks sacrificing the modeling and symbolic capabilities that libraries provide. It also might face new roadblocks as compute hardware continues to evolve. We develop and implement an alternative approach to thermochemistry computation based on symbolic abstractions. From these, performant architecture-specific code can be automatically generated.
@@
The implementation of analytical Jacobians challenges efficiency, and numerical approximations via finite differences require ad hoc tuning of the perturbation size to meet accuracy requirements [15]. Automatic differentiation (AD) is an attractive alternative because it provides the exact derivatives of a function without having to implement them manually in the source code [16]. In practice, AD represents functions as graphs and, as it traverses them, calculates corresponding derivatives via the chain rule. AD libraries and thermochemistry libraries typically belong to disparate compilation units, so the necessary graphs cannot be easily constructed. The approach herein is AD-compatible via the same abstractions that enable accelerated, device-offloaded computation.
@@ 5 Demonstration II: Computation via Integration into Existing Solvers
The preceding demonstrations show a pipeline from symbolic representation to computation, using multiple code transformations and deferred detail addition. Some applications do not benefit from retaining symbolic capabilities, requiring code that instead operates on numeric data immediately within a different control environment than Python. This can be the case for existing HPC flow solvers. Here, we demonstrate that our approach satisfies this. Pyrometheus can generate code that is portable to existing solvers from the symbolic representation without executing the entire pipeline of section 4.
@@ 5.1 Description of the MFC Flow Solver
The MFC algorithm stores and uses coalesced memory arrays for stencil reconstructions. This strategy leads to the relatively high arithmetic intensity and performant code on GPU accelerators. It employs a directional splitting approach that evaluates these quantities once per time step, then reuses them [41, 42, 38]. MFC uses established computational methods, including high-order accurate interface reconstruction, approximate Riemann solvers, and TVD Runge–Kutta time steppers. MFC handles parallel processing via MPI3 and domain decomposition.

### v2

@@ 1.1 Computational Challenges of Combustion Thermochemistry
The evaluation of the thermochemistry often adds significantly to the cost of simulations of reacting flows. This cost primarily follows three sources. The first is an increase in the number of transported variables with respect to simulations of inert flow; the mixture composition can be described by up to hundreds of reacting scalars. The second source is the evaluation of up to thousands of exponentials that appear in expressions representing reaction rates. Last is a time-step restriction that increases the time of solution and stems from the imposed chemical time scale, over 1000 times smaller than that of the flow time scale [3]. These effects compound, and reacting flow simulations can be prohibitively expensive. Compute accelerators (XPUs), including graphic processing units (GPUs) and accelerated processing units (APUs, or superchips), can mitigate this cost.
Our approach complements established thermochemistry software, such as dedicated libraries like Cantera [4], Chemkin [5], and TChem [6], which provide broad modeling capabilities but are limited to CPU hardware. While serving the community well, these are not easily optimized for high-performance simulation, especially on accelerator hardware. This is because the compiled libraries are standalone and live in a different executable than the flow solver. Because of their breadth and structure, adapting such libraries to GPUs is an extensive task that risks conflicting with the goals and success of these tools. Writing new thermochemistry software exclusively for large-scale flow simulations on accelerators risks sacrificing the modeling and symbolic capabilities that libraries provide. It also might face new roadblocks as compute hardware continues to evolve. We develop and implement an alternative approach to thermochemistry computation based on symbolic abstractions. From these, performant architecture-specific code is generated.
@@
The implementation of analytical Jacobians challenges efficiency, and numerical approximations via finite differences require ad hoc tuning of the perturbation size to meet accuracy requirements [15]. Automatic differentiation (AD) is an attractive alternative because it provides the exact derivatives of a function without having to implement them manually in the source code [16]. In practice, AD represents functions as graphs and calculates corresponding derivatives via the chain rule as it traverses them. AD libraries and thermochemistry libraries typically belong to disparate compilation units, so the necessary graphs cannot be easily constructed. The approach herein is AD-compatible via the same abstractions that enable accelerated, device-offloaded computation.
@@ 5 Demonstration II: Computation via Integration into Existing Solvers
The preceding demonstrations show a pipeline from symbolic representation to computation, using multiple code transformations and deferred detail addition. Some applications do not benefit from retaining symbolic capabilities, requiring code that instead operates on numeric data immediately within a different control environment than Python. This situation is common when using existing flow solvers. Here, we demonstrate that Pyrometheus can generate code that is portable to existing solvers from the symbolic representation without executing the entire pipeline of section 4.
@@ 5.1 Description of the MFC Flow Solver
The MFC algorithm stores and uses coalesced memory arrays for stencil reconstructions. This strategy leads to the relatively high arithmetic intensity and performant code on GPU accelerators. It employs a directional splitting approach that evaluates these quantities once per time step, then reuses them [41, 42, 38]. MFC uses established computational methods, including high-order accurate interface reconstruction, approximate Riemann solvers, and Runge–Kutta time steppers. MFC handles parallel processing via MPI3 and domain decomposition.
@@ 5.4 Pyrometheus-enabled MFC Simulations of Detonations
5.5 Pyrometheus-enabled MFC Simulations of Flame–Vortex Interactions
To demonstrate our capability for simulating thermochemistry and transport, we simulate a two-dimensional premixed hydrogen–air flame perturbed by two superimposed vortices. This simulation closely follows the configuration of Yu et al. [49] and Bando et al. [50]. In contrast with the previous one-dimensional detonation, this case incorporates mixture-averaged mass diffusion fluxes and corresponding transport coefficients 25, 27 and 30. Reactions are modeled using the hydrogen subset of the GRI-3.0 mechanism. Fifth-order accurate WENO reconstruction and an HLLC approximate Riemann solver are used to compute the fluxes, following [35].
[…]
The flame–vortex interaction is shown in fig. 13. We compare the flame shape with the results of Bando et al. [50]. Close agreement in flame shape is observed. Differences in absolute spatial position are due to the lack of sufficient geometric information in the reference solution. Still, the vortex structures observed in our simulation also show good agreement with the results of Yu et al. [49], where similar features in the vorticity field are reported. Vortex structures downstream of the flame arise due to flame curvature, which induces baroclinicity. The agreement in flame and flow structures supports the correctness of the implementation.
As stated, these simulations necessitate the computation of transport coefficients, which are provided to MFC by Pyrometheus. This verification is in contrast with the simulation of one-dimensional detonation of fig. 12. For this hydrogen–air mechanism, the cost of calculating mixture-averaged transport properties is of total run time, which is of the time that it takes to assemble the corresponding mass diffusion fluxes. This cost balance indicates the efficiency of our Pyrometheus implementation for transport.
The simulations presented in this section showcase how Pyrometheus overcomes the challenges associated with coupling exascale flow solvers with computational models for chemical kinetics, thermodynamics, and transport. In contrast with pre-compiled libraries, Pyrometheus places the computation of the thermochemistry and transport at the same compilation level as the flow solver. MFC then offloads the fully coupled flow–combustion computation to GPUs using compiler directives (at odds with pre-compiled libraries), achieving near roofline performance (fig. 11).